OpenStudy (anonymous):

sqrt(4x+7)² = sqrt(4x+7) * sqrt(4x+7) right?

OpenStudy (anonymous):

\[(\sqrt{4x+7}_{})^{2} = (\sqrt{4x+7})(\sqrt{4x+7})\]

OpenStudy (anonymous):

Can someone show me how I go from \[(\sqrt{4x+7})(\sqrt{4x+7}) \to 4x + 7\]

OpenStudy (saifoo.khan):

oh man! \[(\sqrt(4x+7))^2 = (4x+7)^\frac22\]

OpenStudy (anonymous):

\[(\sqrt{4x+7})^{2} = ((4x+7)_{2}^{1})_{}^{2}\] Then power of a power law means that \[(a ^{m})^{n} = a ^{m*n}\] So the 1/2 exponent * the 2 exponents equals 2/2 which is 1 right?

OpenStudy (saifoo.khan):

Yes sir.